OpenStudy (anonymous):

Six is the geometric mean of 18 and what number? 4 3 2 9

OpenStudy (dariyana):

2

OpenStudy (anonymous):

itsss 2

OpenStudy (anonymous):

Let other number be : b \[G.M = \sqrt{a \times b}\] G.M. = 6 a = 18 Plug in the values: \[6 = \sqrt{18 \times b}\] \[18 \times b = 36\] \[b = 2\]